Five times the first of three consecutive even integers is fourteen more than three times the second. What is the third consecutive integer?: Five times the first of three consecutive even integers is fourteen more than three times the second. What is the third consecutive integer?
Answer by nerdybill(1129)

Five times the first of three consecutive even integers is fourteen more than three times the second. What is the third consecutive integer?
.
Let x = 1st of three consecutive integers
x+2 = 2nd of three consecutive integers
x+4 = 3rd of three consecutive integers
.
From:"Five times the first of three consecutive even integers is fourteen more than three times the second." we get:
5x = 3(x+2)+14
.
Solving for x:
5x = 3(x+2)+14
5x = 3x+6+14
5x = 3x+20
2x = 20
x = 10
.
The "3rd of the three consecutive integers" is:
x+4 = 10+4 = 14